# Perfect quantum state transfer through a chaotic spin chain via many-body scars

### v1

@@ I Introduction
A central question is whether QMBS can be leveraged to reliably process quantum information. Prior studies have primarily focused on finding new models hosting QMBS [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25], characterising their properties [26, 27, 28, 29, 30, 31, 32] or amplifying the oscillations of local observables [33, 34, 35]. While these features suggest the potential for coherent information storage and manipulation, demonstrations that exploit QMBS for actual quantum information tasks are rare (exceptions include Refs. [36, 37], in the context of quantum sensing and metrology).
Figure 1: (a) The aim of the quantum state transfer protocol is to transmit an arbitrary spin-1/2 state $|\psi\rangle=\alpha|0\rangle+\beta|1\rangle$ across the chain. In this paper we focus on state transfer implemented by time-independent Hamiltonian dynamics. (b) Perfect state transfer via $\hat{H}_{\rm PST}$ can be understood intuitively as precession of an effective spin-$(N-1)/2$, embedded in the many-body Hilbert space, which transports a spin excitation across the chain. This effective spin is preserved in our many-body scarred Hamiltonian $\hat{H}_{\rm scar}=\hat{H}_{\rm PST}+\sum_{n}\hat{P}_{n}\hat{h}_{n}\hat{P}_{n}$, while the remainder of the spectrum is thermalised.
@@
Our results demonstrate that perfect quantum state transfer can occur in a many-body system that is quantum chaotic, challenging the conventional wisdom. In other words, by harnessing the special structure of QMBS, quantum chaos need not be an obstacle to control. This establishes a novel route to exploiting weak ergodicity breaking in chaotic systems, and opens new directions for the design of quantum technologies in settings that do not rely on either integrability or many-body localisation.
@@ II Perfect state transfer in a spin-1/2 chain
To intuitively understand this perfect state transfer, let us first observe that the Hamiltonian has a $U(1)$ symmetry, corresponding to the total magnetisation $\hat{M}=\sum_{n}\hat{Z}_{n}$ being a conserved quantity $[\hat{H}_{\rm PST},\hat{M}]=0$. The initial state superposition in Eq. 2 has two components. First, the $|0\rangle^{\otimes N}$ component is an eigenstate in the subspace of maximal total magnetisation $M=N$, and therefore only accumulates a phase $|0\rangle^{\otimes N}\to e^{-itN\omega/2}|0\rangle^{\otimes N}$ during the evolution. Second, the $|1\rangle|0\rangle^{\otimes(N-1)}$ is an element of the next lowest subspace $M=N-2$, and therefore evolves entirely within this subspace, which is spanned by the states:
@@
It is straightforward to verify that the Hamiltonian restricted to this subspace is (up to an added term proportional to the identity in the subspace) equal to:
@@
where $\hat{J}^{+}\equiv\sum_{n=1}^{N-1}\sqrt{n(N-n)}|n+1\rangle\langle n|$ , $\hat{J}^{x}=\frac{1}{2}(\hat{J}^{+}+\hat{J}^{-})$, $\hat{J}^{y}=\frac{i}{2}(\hat{J}^{-}-\hat{J}^{+})$ and $\hat{J}^{z}=\frac{1}{2}[\hat{J}^{+},\hat{J}^{-}]=\frac{1}{2}\sum_{n=}^{N}(2n-N-1)|n\rangle\langle n|$. These $\hat{J}^{\mu}$ are the familiar spin angular momentum operators for a large spin-$J$ particle, with $J=(N-1)/2$. Starting from the lowest weight spin state in the $z$-direction $|n=1\rangle=|10...0\rangle$, evolution by $\hat{\mathsf{H}}_{\rm subspace}=2\lambda\hat{J}^{x}$ causes the effective spin-$J$ to precess around its $x$-axis to the highest weight spin state $|n=N\rangle=|0...01\rangle$ after an evolution time $t=\tau_{1}=\pi/2\lambda$, implementing the transfer of the excitation across the chain [see Fig. 1(b)]. Moreover, we expect perfect state transfer at periodic intervals thereafter, at the times $t=\tau_{m}=\pi(m-1/2)/\lambda$ for $m=1,2,3,\ldots$, as the effective spin continues to precess.
@@
Following the same line of reasoning as above, the $\beta$-component $|1\rangle^{\otimes N}$ accummulates a phase, while the $\alpha$-component $|0\rangle|1\rangle^{\otimes(N-1)}$ evolves as a precessing spin-$(N-1)/2$ particle in a subspace spanned by the states $|\overline{n}\rangle=|1_{1}\ldots 1_{n-1}0_{n}1_{n+1}\ldots 1_{N}\rangle$. For this initial state, however, the final unitary correction of the last spin should be modified $\hat{V}(\tau_{1})\to\hat{V}^{\dagger}(\tau_{1})$, and the transfer fidelity is given by $\overline{F}(t)=\langle\psi|\hat{V}(t)\hat{\rho}_{N}(t)\hat{V}^{\dagger}(t)|\psi\rangle$.
@@ IV Perfect state transfer in a chaotic spin-1/2 chain with QMBS
Just like for the Hamiltonian $\hat{H}_{\rm thermal}$ condidered previously, the second term $\sum_{n}\hat{P}_{n}\hat{h}_{n}\hat{P}_{n}$ generally breaks the integrability of the model. For example, let us again choose $\hat{h}_{n}=\hat{\mathbb{I}}_{2}^{\otimes(n-2)}\otimes\hat{h}\otimes\hat{\mathbb{I}}_{2}^{\otimes(N-n-1)}$ where $\hat{h}$ is a randomly generated Hermitian operator acting on three spins. In Fig. 2(c) we show that the level spacing statistics of $\hat{H}_{\rm scar}$ for this example indicate that the Hamiltonian is quantum chaotic. At first glance, one might expect that this will prevent state transfer through the spin chain. However, the crucial difference between $\hat{H}_{\rm thermal}$ the $\hat{H}_{\rm scar}$ is that the latter Hamiltonian hosts QMBS, due to its special structure involving the local projectors $\hat{P}_{n}$. This is evident in the eigenstate entanglement entropies shown in Fig. 2(e): the vast majority of eigenstates are thermal (they have a large entanglement) but there is a small subset of $N+1$ non-thermal eigenstates near the middle of the energy spectrum with significantly lower entanglement. These are the QMBS eigenstates.
@@
It is straightforward to show mathematically that the Hamiltonian $\hat{H}_{\rm scar}$ has a set of $N+1$ QMBS. First, we observe that the state $|0\rangle^{\otimes N}$, which is an eigenstate of the first term $\hat{H}_{\rm PST}|0\rangle^{\otimes N}=\frac{\omega N}{2}|0\rangle^{\otimes N}$, is annihilated by every projector $\hat{P}_{n}$, and therefore also by their sum $\sum_{n}\hat{P}_{n}\hat{h}_{n}\hat{P}_{n}|0\rangle^{\otimes N}=0$. It follows that this state is a QMBS of the total Hamiltonian $\hat{H}_{\rm scar}|0\rangle^{\otimes N}=\frac{\omega N}{2}|0\rangle^{\otimes N}$. It is clearly visible as a zero-entanglement outlier in the eigenstate entanglement entropy plot in Fig. 2(e).
@@
Next, we observe that the states $|n\rangle$ (defined in Eq. 5) are annihilated by every projector $\hat{P}_{n}$, i.e., $\hat{P}_{n}|n^{\prime}\rangle=0$ for all $n$, $n^{\prime}$. The sum of projectors therefore also annihilates these states, $\sum_{n}\hat{P}_{n}\hat{h}_{n}\hat{P}_{n}|n^{\prime}\rangle=0$, and the subspace spanned by $\{|n\rangle\}_{n=1}^{N}$ is “invisible” to the Hamiltonian perturbation. However, we already know that the Hamiltonian $\hat{H}_{\rm PST}$ restricted to this subspace is given by $\hat{\mathsf{H}}_{\rm subspace}=2\lambda\hat{J}^{x}$ (Eq. 6). Let $|x_{n}\rangle=e^{i\pi\hat{J}^{y}/2}|n\rangle$ be the eigenstates of $\hat{J}^{x}$, with the harmonically spaced eigenvalues $\hat{J}^{x}|x_{n}\rangle=\frac{1}{2}(2n-N-1)|x_{n}\rangle$. It is clear that these must be eigenstates of the total Hamiltonian, with $\hat{H}_{\rm scar}|x_{n}\rangle=[\omega(N-2)/2+\lambda(2n-N-1)]|x_{n}\rangle$ [the $\omega(N-2)/2$ component comes from the total magnetization in the $M=N-2$ subspace]. Since the $|x_{n}\rangle$ are a superposition of at most $N$ product states $|n\rangle$, their half-chain entanglement entropy is bounded by the sub-volume scaling $S_{\rm ent}\leq\log\frac{N}{2}$, i.e., they are non-thermal QMBS. These are clearly visible as the band of low (but non-zero) entanglement outliers in Fig. 2(e).
@@
We emphasise that the perfect state transfer generated by $\hat{H}_{\rm scar}$ does not rely on any local conserved quantities of the Hamiltonian, such as total magnetisation. In fact, the $U(1)$ symmetry present in $\hat{H}_{\rm PST}$ is broken by the added $\sum_{n}\hat{P}_{n}\hat{h}_{n}\hat{P}_{n}$ term in the Hamiltonian, and $\hat{H}_{\rm scar}$ has no obvious local conserved quantities (apart from the trivial identity operator and the Hamiltonian itself). However, the Hamiltonian does have a non-local symmetry, $[\hat{H}_{\rm scar},\hat{J}^{x}]=0$, where $\hat{J}^{x}$ is the effective spin operator defined below Eq. 6, which protects the dynamics necessary for perfect state transfer.
@@ V Perfect state transfer in a chaotic spin-1 chain with QMBS
Finally, we observe that for spin-1 particles, any two-body interaction $\hat{h}$ can be decomposed as a linear combination of $813^{2}\times 3^{2}=$ 81 linearly independent Hermitian operators (i.e, 81 different interaction types). The effect of the projectors in $\hat{P}_{n,n^{\prime}}\hat{h}_{n,n^{\prime}}\hat{P}_{n,n^{\prime}}$, with $\hat{P}_{n,n^{\prime}}$ given by Eq. 16, is to annihilate 45 of the possible interaction types, while commuting with the remaining 36 interaction types. So, our state transfer protocol is robust to 36 of the possible 81 kinds of two-spin interactions. An important direction of future work might be to combine this partial robustness with other error mitigation techniques. For example, using dynamical decoupling, it may be possible to drive the system away from the interaction types against which the protocol is fragile and towards the interaction types against which the protocol is robust. The pulse design requirements are likely to be less severe than, for example, if all interaction types were required to be suppressed [53].
@@ VI Discussion
We have demonstrated that perfect quantum state transfer is possible through a spin chain even in the presence of quantum chaos, provided the system Hamiltonian hosts QMBS. The result is striking, as chaotic dynamics are typically associated with rapid delocalisation and information scrambling, which would intuitively seem to hinder coherent state transfer. The presence of QMBS enables non-ergodic dynamics from a small subset of initial conditions, within an otherwise chaotic system, thereby supporting efficient state transfer.
@@
We note that our scarred models feature an embedded subspace with an $SU(2)$ structure that protects the coherent dynamics underlying the state transfer. Several previously known scarred models, including the PXP model, also exhibit an exact or approximate $SU(2)$ structure [33, 28]. It would be interesting to explore whether other $SU(2)$-based QMBS systems could similarly support quantum information tasks such as state transfer. Finally, while QMBS can protect against certain perturbations, they may themselves be fragile under more generic interactions. Enhancing their robustness, perhaps through dynamical decoupling or other control techniques, remains an important open challenge.
@@ Acknowledgements.
This publication has emanated from research conducted with the financial support of Taighde Éireann – Research Ireland under Grant number 22/PATH-S/10812.

### v2

@@ I Introduction
A central question is whether QMBS can be leveraged to reliably process quantum information. Prior studies have primarily focused on finding new models hosting QMBS [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27], characterising their properties [28, 29, 30, 31, 32, 33, 34] or amplifying the oscillations of local observables [35, 36, 37]. While these features suggest the potential for coherent information storage and manipulation, demonstrations that exploit QMBS for actual quantum information tasks are rare (exceptions include Refs. [38, 39], in the context of quantum sensing and metrology).
Figure 1: (a) The aim of the quantum state transfer protocol is to transmit an arbitrary spin-1/2 state $|\psi\rangle=\alpha|0\rangle+\beta|1\rangle$ across the chain. In this paper we focus on state transfer implemented by time-independent Hamiltonian dynamics. (b) Perfect state transfer via $\hat{H}_{\rm PST}$ can be understood intuitively as precession of an effective spin-(N1)/2(N-1)/2, embedded in the many-body Hilbert space, which transports a spin excitation across the chain. This effective spin is preserved in our many-body scarred Hamiltonian $\hat{H}_{\rm scar}=\hat{H}_{\rm PST}+\sum_{n}\hat{P}_{n}\hat{h}_{n}\hat{P}_{n}$, while the remainder of the spectrum is thermalised.
@@
Our results demonstrate that perfect quantum state transfer can occur in a many-body system that is quantum chaotic, challenging the conventional wisdom. In other words, by harnessing the special structure of QMBS, quantum chaos need not be an obstacle to control.
@@ II Perfect state transfer in a spin-1/2 chain
To intuitively understand this perfect state transfer, let us first observe that the Hamiltonian has a $U(1)$ symmetry, corresponding to the total magnetisation $\hat{M}=\sum_{n}\hat{Z}_{n}$ being a conserved quantity $\hat{H}_{\rm PST},\hat{M}]=0$. The initial state superposition in Eq. 2 has two components. First, the $|0\rangle^{\otimes N}$ component is an eigenstate in the subspace of maximal total magnetisation $M=N$, and therefore only accumulates a phase $|0\rangle^{\otimes N}\to e^{-itN\omega/2}|0\rangle^{\otimes N}$ during the evolution. Second, the $|1\rangle|0\rangle^{\otimes(N-1)}$ is an element of the next lowest subspace M=N2M=N-2, and therefore evolves entirely within this subspace, which is spanned by the states:
@@
It is straightforward to verify that the Hamiltonian $\hat{H}_{\rm PST}$ restricted to this subspace is (up to an added term proportional to the identity in the subspace) equal to:
@@
where $\hat{J}^{+}\equiv\sum_{n=1}^{N-1}\sqrt{n(N-n)}|n+1\rangle\langle n|$ , $\hat{J}^{x}=\frac{1}{2}(\hat{J}^{+}+\hat{J}^{-})$, $\hat{J}^{y}=\frac{i}{2}(\hat{J}^{-}-\hat{J}^{+})$ and $\hat{J}^{z}=\frac{1}{2}[\hat{J}^{+},\hat{J}^{-}]=\frac{1}{2}\sum_{n=}^{N}(2n-N-1)|n\rangle\langle n|$. These $\hat{J}^{\mu}$ are the familiar spin angular momentum operators for a large spin-$J$ particle, with J=(N1)/2J=(N-1)/2. Starting from the lowest weight spin state in the $z$-direction $|n=1\rangle=|10...0\rangle$, evolution by $\hat{\mathsf{H}}_{\rm subspace}=2\lambda\hat{J}^{x}$ causes the effective spin-$J$ to precess around its $x$-axis to the highest weight spin state $|n=N\rangle=|0...01\rangle$ after an evolution time $t=\tau_{1}=\pi/2\lambda$, implementing the transfer of the excitation across the chain [see Fig. 1(b)]. Moreover, we expect perfect state transfer at periodic intervals thereafter, at the times $t=\tau_{m}=\pi(m-1/2)/\lambda$ for $m=1,2,3,\ldots$, as the effective spin continues to precess.
@@
Following the same line of reasoning as above, the $\beta$-component $|1\rangle^{\otimes N}$ accummulates a phase, while the $\alpha$-component $|0\rangle|1\rangle^{\otimes(N-1)}$ evolves as a precessing spin-(N1)/2(N-1)/2 particle in a subspace spanned by the states $|\overline{n}\rangle=|1_{1}\ldots 1_{n-1}0_{n}1_{n+1}\ldots 1_{N}\rangle$. For this initial state, however, the final unitary correction of the last spin should be modified $\hat{V}(\tau_{1})\to\hat{V}^{\dagger}(\tau_{1})$, and the transfer fidelity is given by $\overline{F}(t)=\langle\psi|\hat{V}(t)\hat{\rho}_{N}(t)\hat{V}^{\dagger}(t)|\psi\rangle$.
@@ IV Perfect state transfer in a chaotic spin-1/2 chain with QMBS
Just like for the Hamiltonian $\hat{H}_{\rm thermal}$ condidered previously, the second term $\sum_{n}\hat{P}_{n}\hat{h}_{n}\hat{P}_{n}$ generally breaks the integrability of the model. For example, let us again choose $\hat{h}_{n}=\hat{\mathbb{I}}_{2}^{\otimes(n-2)}\otimes\hat{h}\otimes\hat{\mathbb{I}}_{2}^{\otimes(N-n-1)}$ where $\hat{h}$ is a randomly generated Hermitian operator acting on three spins. In Fig. 2(c) the level spacing statistics of $\hat{H}_{\rm scar}$ for this example indicate that the Hamiltonian is quantum chaotic. At first glance, one might expect that this will prevent state transfer through the spin chain. However, the crucial difference between $\hat{H}_{\rm thermal}$ the $\hat{H}_{\rm scar}$ is that the latter Hamiltonian hosts QMBS, due to its special structure involving the local projectors $\hat{P}_{n}$. This is evident in the eigenstate entanglement entropies shown in Fig. 2(e): the vast majority of eigenstates are thermal (they have a large entanglement) but there is a small subset of $N+1$ non-thermal eigenstates near the middle of the energy spectrum with significantly lower entanglement. These are the QMBS eigenstates.
@@
It is straightforward to show mathematically that the Hamiltonian $\hat{H}_{\rm scar}$ has a set of $N+1$ QMBS. First, we observe that the state $|0\rangle^{\otimes N}$, which is an eigenstate of the first term $\hat{H}_{\rm PST}|0\rangle^{\otimes N}=\frac{\omega N}{2}|0\rangle^{\otimes N}$, is annihilated by every projector $\hat{P}_{n}$, and therefore also by the sum $\sum_{n}\hat{P}_{n}\hat{h}_{n}\hat{P}_{n}|0\rangle^{\otimes N}=0$. It follows that this state is a QMBS of the total Hamiltonian $\hat{H}_{\rm scar}|0\rangle^{\otimes N}=\frac{\omega N}{2}|0\rangle^{\otimes N}$. It is clearly visible as a zero-entanglement outlier in the eigenstate entanglement entropy plot in Fig. 2(e).
@@
Next, we observe that the states $|n\rangle$ (defined in Eq. 5) are annihilated by every projector $\hat{P}_{n}$, i.e., $\hat{P}_{n}|n^{\prime}\rangle=0$ for all $n$, $n^{\prime}$. The second term in Eq. 10 therefore also annihilates these states, $\sum_{n}\hat{P}_{n}\hat{h}_{n}\hat{P}_{n}|n^{\prime}\rangle=0$, and the subspace spanned by $\{|n\rangle\}_{n=1}^{N}$ is “invisible” to the Hamiltonian perturbation. However, we already know that the Hamiltonian $\hat{H}_{\rm PST}$ restricted to this subspace is given by $\hat{\mathsf{H}}_{\rm subspace}=2\lambda\hat{J}^{x}$ (Eq. 6). Let $|x_{n}\rangle=e^{i\pi\hat{J}^{y}/2}|n\rangle$ be the eigenstates of $\hat{J}^{x}$, with the harmonically spaced eigenvalues $\hat{J}^{x}|x_{n}\rangle=\frac{1}{2}(2n-N-1)|x_{n}\rangle$. It is clear that these must be eigenstates of the total Hamiltonian, with $\hat{H}_{\rm scar}|x_{n}\rangle=[\omega(N-2)/2+\lambda(2n-N-1)]|x_{n}\rangle$ [the $\omega(N-2)/2$ component comes from the total magnetization in the M=N2M=N-2 subspace]. Since the $|x_{n}\rangle$ are a superposition of at most $N$ product states $|n\rangle$, their half-chain entanglement entropy is bounded by the sub-volume scaling $S_{\rm ent}\leq\log\frac{N}{2}$, i.e., they are non-thermal QMBS. These are clearly visible as the band of low (but non-zero) entanglement outliers in Fig. 2(e).
@@
We emphasise that the perfect state transfer generated by $\hat{H}_{\rm scar}$ does not rely on any local conserved quantities of the Hamiltonian, such as total magnetisation. In fact, the $U(1)$ symmetry present in $\hat{H}_{\rm PST}$ is broken by the added $\sum_{n}\hat{P}_{n}\hat{h}_{n}\hat{P}_{n}$ term in the Hamiltonian, and $\hat{H}_{\rm scar}$ has no obvious local conserved quantities (apart from the trivial identity operator and the Hamiltonian itself). However, the Hamiltonian does have a non-local symmetry, $\hat{H}_{\rm scar},\hat{J}^{x}]=0$, where $\hat{J}^{x}$ is the effective spin operator defined below Eq. 6, which protects the dynamics necessary for perfect state transfer.
@@ V Perfect state transfer in a chaotic spin-1 chain with QMBS
Finally, we observe that for spin-1 particles, any two-body interaction $\hat{h}$ can be decomposed as a linear combination of $3^{2}\times 3^{2}=81$ linearly independent Hermitian operators (i.e, 81 different interaction types). The effect of the projectors in $\hat{P}_{n,n^{\prime}}\hat{h}_{n,n^{\prime}}\hat{P}_{n,n^{\prime}}$, with $\hat{P}_{n,n^{\prime}}$ given by Eq. 16, is to annihilate 45 of the possible interaction types, while commuting with the remaining 36 interaction types. So, our state transfer protocol is robust to 36 of the possible 81 kinds of two-spin interactions. An important direction of future work might be to combine this partial robustness with other error mitigation techniques [55]. For example, using dynamical decoupling, it may be possible to drive the system away from the interaction types against which the protocol is fragile and towards the interaction types against which the protocol is robust. The pulse design requirements are likely to be less severe than, for example, if all interaction types were required to be suppressed [56].
@@ VI Discussion
We have demonstrated that perfect quantum state transfer is possible through a spin chain even in the presence of quantum chaos, in a particular set of Hamiltonians hosting QMBS. The result is striking, as chaotic dynamics are typically associated with rapid delocalisation and information scrambling, which would intuitively seem to hinder coherent state transfer. The presence of QMBS enables non-ergodic dynamics from a small subset of initial conditions, within an otherwise chaotic system, thereby supporting perfect state transfer.
@@
Our scarred Hamiltonians in Eqs. 10 and 15 bear a strong resemblance to the class of scarred Hamiltonians introduced by Shiraishi and Mori [3]. Despite these similarities, in Appendix A we show that our Hamiltonians do not fall exactly into this class. The Shiraishi-Mori construction could be used to embed the integrable spin-1/2 state transfer Hamiltonian in a larger chaotic spin-1 chain in a relatively trivial manner that differs from our approach. However, in Appendix B we show that this would give an unnecessarily large scar subspace, while our scarred Hamiltonian is “minimal”: it has exactly the scar subspace required for perfect state transfer, without any superfluous scar degrees of freedom, which allows us to find a larger set of interactions to which the state transfer protocol is robust.
We note that our scarred models feature an embedded subspace with an $SU(2)$ structure that protects the coherent dynamics underlying the state transfer. Several previously known scarred models, including the PXP model, also exhibit an exact or approximate $SU(2)$ structure [35, 30]. It would be interesting to explore whether other $SU(2)$-based QMBS systems could similarly support quantum information tasks such as state transfer, perhaps through solitonic excitations [57]. Also, while QMBS can protect against certain perturbations, they may themselves be fragile under more generic interactions (see Appendix C). Enhancing their robustness, perhaps through dynamical decoupling or other control techniques, remains an important open challenge.
@@ Acknowledgements.
This publication has emanated from research conducted with the financial support of Taighde Éireann – Research Ireland under Grant number 22/PATH-S/10812. The data that support the findings of this article are openly available at [58].
Appendix A Relationship of our models to the Shiraishi-Mori projector embedding construction
The construction of our scarred Hamiltonians $\hat{H}_{\rm scar}$ (Eq. 10) and $\hat{H}_{\rm scar}^{(1)}$ (Eq. 15) is similar to, and inspired by, the projector-embedding method, due to Shiraishi and Mori [3]. However, our model does not fit exactly into their class of models.
The Shiraishi-Mori method builds a local Hamiltonian of the form $\hat{H}=\hat{H}^{\prime}+\sum_{n}\hat{P}_{n}\hat{h}_{n}\hat{P}_{n}$ where $\hat{H}^{\prime},\hat{P}_{n}]=0$ for all $n$. The scar subspace is the set of states $\mathcal{S}=\{|\Psi\rangle:\hat{P}_{n}|\Psi\rangle=0,\forall n\}$ that are simultaneously annihilated by all local projectors $\hat{P}_{n}$. Our model has the similar form $\hat{H}=\hat{H}_{\rm PST}+\sum_{n}\hat{P}_{n}\hat{h}_{n}\hat{P}_{n}$, but without the commutation relation being satisfied, i.e., $\hat{H}_{\rm PST},\hat{P}_{n}]\neq 0$. Instead, our projectors commute with the non-local operator $\hat{\mathsf{H}}_{\rm subspace}=2\lambda\hat{J}^{x}$, which is the Hamiltonian $\hat{H}_{\rm PST}$ restricted to its subspace of total magnetization M=N2M=N-2 (i.e., we have $\hat{\mathsf{H}}_{\rm subspace},\hat{P}_{n}]=0$ for all $n$). Also, our scar subspace is not $\mathcal{S}$, rather it is the subspace of $\mathcal{S}$ in the M=N2M=N-2 and $M=N$ magnetisation sectors. So, although our construction has similarities to the Shiraish-Mori projector embedding method, there are also several differences.
Appendix B “Trivial” embedding of spin-1/2 state transfer Hamiltonian in a spin-1 chain
It would be possible to embed the integrable spin-1/2 state transfer Hamiltonian $\hat{H}_{\rm PST}$ in a larger chaotic spin-1 chain in a relatively trivial manner by the Shiraishi-Mori projector embedding method. In our spin-1 example of Sec. V, this would be achieved by replacing the projector:
[…]
(given in our Eq. 16) with the projector:
[…]
However, this replacement would give a scar subspace $\mathcal{S}={\rm span}\{|i_{1}i_{2}\ldots i_{N}\rangle:i_{j}\in\pm\}$, which is $2^{N}$-dimensional, when only an $(N+1)$-dimensional subspace is required to implement perfect state transfer. In this sense, the “trivial” construction would provide an unnecessarily large scar subspace, while our embedding is “minimal”: it gives exactly the required scar subspace for perfect state transfer, without any superfluous scar degrees of freedom. Moreover, the trivial embedding method would not work at all in our spin-1/2 chain example in Sec. IV, since then the $2^{N}$-dimensional scar subspace would be the entire $2^{N}$-dimensional Hilbert space of the chain. An advantage of our “minimal” scar subspace embedding is that allows us to identify a larger set of local interaction terms ($\hat{P}_{n}\hat{h}_{n}\hat{P}_{n}$ in our notation) to which our protocol is robust, compared to the trivial embedding. For example, state transfer through our spin-1 chain, using the local projectors $\hat{P}$ is explicitly robust to local interactions of the form:
$\hat{h}$ $=$ $a|-,-\rangle\langle 0,+|+b|-,-\rangle\langle+,0|+c|-,-\rangle\langle 0,-|$
$\qquad+d|-,-\rangle\langle-,0|+f|-,-\rangle\langle 0,0|+{\rm h.c.},$
since for this form of $\hat{h}$ we have $\hat{P}\hat{h}\hat{P}=\hat{h}$. However, $\hat{P}^{(\rm triv)}\hat{h}\hat{P}^{(\rm triv)}\neq\hat{h}$ for the corresponding local projectors $\hat{P}^{(\rm triv)}$ of the trivial embedding, making the robustness of state transfer to interactions of type $\hat{h}$ difficult to see for the trivial embedding.
Appendix C State transfer fidelity in the presence of non-projected perturbations
As discussed at the end of Sec. V, perfect state transfer through the spin-1 chain is robust to 36 of the possible 81 types of two-spin interaction. In this section we examine the transfer fidelity as a function of perturbation strength, for perturbations that can destroy the QMBS.
We consider our spin-1 chain model $\hat{H}_{\rm scar}^{(1)}$ (given in Eq. 15 of our manuscript) and three different perturbations: $\hat{H}_{\rm local-X}=\epsilon\hat{S}^{x}_{N/2}$, $\hat{H}_{\rm global-X}=\epsilon\sum_{n}\hat{S}^{x}_{n}$, and $\hat{H}_{\rm global-YY}=\epsilon\sum_{n}\hat{S}^{y}_{n}\hat{S}^{y}_{n+1}$ where:
$\hat{S}^{x}=\left(\begin{array}[]{ccc}0&1&0\\ 1&0&1\\ 0&1&0\end{array}\right),\quad\hat{S}^{y}=\left(\begin{array}[]{ccc}0&-i&0\\ i&0&-i\\ 0&i&0\end{array}\right),$ (19)
in the spin-1 basis $\{|-\rangle,|0\rangle,|+\rangle\}$. For each perturbation type, we numerically compute the state transfer infidelity 1F(t)1-F(t) at the time $t=\tau_{1}=\pi/2\lambda$ (which is the time of perfect state transfer when $\epsilon=0$). The blue lines in Figs. 4(a,b,c) show that the infidelity scales as $1-F(\tau_{1})\sim\epsilon^{2}$ with the perturbation strength $\epsilon$, for all perturbation types. For comparison, we also plot the infidelity as a function of $\epsilon$ when the perturbation is added to the integrable perfect state transfer Hamiltonian $\hat{H}_{\rm PST}^{(1)}$ (given in Eq. 12 of our manuscript). The orange lines in Figs. 4(a,b,c) show the scaling $1-F(\tau_{1})\sim\epsilon^{2}$ for all perturbation types in this case too. Interestingly, the state transfer fidelity is worse when the perturbation is added to the integrable Hamiltonian than when it is added to the non-integrable scarred Hamiltonian. The reason for this is unclear and deserves further study.
Figure 4: The state transfer infidelity, at time $t=\pi/2\lambda$, in the presence of three different perturbation types: (a) $\hat{H}_{\rm global-X}=\epsilon\sum_{n}\hat{S}_{n}^{x}$, (b) $\hat{H}_{\rm local-X}=\epsilon\hat{S}_{N/2}^{x}$, (c) $\hat{H}_{\rm global-YY}=\epsilon\sum_{n}\hat{S}_{n}^{y}\hat{S}_{n+1}^{y}$. Here, $N=8$ and all other initial conditions and parameters in $\hat{H}_{\rm scar}^{(1)}$ are generated in the same way as in Fig. 3.